OpenStudy (anonymous):

What is 4.32 × 10^3 ? A. 432 B. 4,320 C. 43,200 D. 432,000 This is D. Right?

OpenStudy (anonymous):

4.32*10*10*10 =43.2*10*10=432*10=4320

OpenStudy (anonymous):

so its B

OpenStudy (anonymous):

its simple,dont be confused with easy multiplication

OpenStudy (jadeishere):

Want to know a trick @Teehee321 ? It's easy to use in scientific notation

OpenStudy (anonymous):

But I thought it would be be D because 4.32 is already there then you add 3 numbers so it would be 4,32000

OpenStudy (anonymous):

o_o

OpenStudy (anonymous):

thats 4 comma 32000 which is 4 lakh 32 thousand 4 point 32 in equation is 4 point 32 which is a number btwn 4 and 5

OpenStudy (jadeishere):

Nononononono, there's the first digit, which is four in this case then the next three digits will be used so there's 4 then there's the 3 and 2 that takes away two of the three digits so there's 432, then add another 0 4320

OpenStudy (anonymous):



OpenStudy (anonymous):

4,320

OpenStudy (anonymous):